OpenStudy (anonymous):

Write the standard form of the equation of the line through the pair of points (3, -7) and (6, 5). Write the equation using only integer coefficients.

OpenStudy (anonymous):

slope of line,m = (y2-y1)/(x2-x1) std eq. of line y-y1 = m(x-x1)

OpenStudy (anonymous):

yeah I know that part